# Trainable Dynamic Mask Sparse Attention

### v1

@@ 4 Method
$h^{\prime}_{t}$ $=\operatorname{concat}(o_{1},o_{2},\ldots,o_{n_{h}})W^{OD}$
$where\quad a_{t}\in\mathbb{R}^{n_{h}\times t}\quad o_{t}\in\mathbb{R}^{n_{h}\times d_{h}}\quad W^{OD}\in\mathbb{R}^{n_{h}\times d_{h}\times d_{model}}\quad h^{\prime}_{t}\in\mathbb{R}^{d_{model}}$
Mask Structure and Weights Structure.
The design of our two core components draws inspiration from existing efficient sequence transformation architectures, but innovates in terms of content-aware sparsity and position-aware sparsity.
Our dynamic mask generation mechanism draws inspiration from structured State-Space [13]. While State-Space achieve selectivity by compressing historical information into a fixed-size hidden state through input-dependent state transitions, DMA takes a fundamentally different approach by retaining the complete historical information while implementing selectivity through content-aware attention masks. The zero-order hold method ensures that the dynamic weights remain stable across sequence positions while providing learnable time-scale parameters to control the selection process, and modulating the strength of the selection signal across different attention heads, bridging the gap between the recurrent nature of State-Space and the parallel computation advantages of Self-Attention.
Our sparse attention weights computation mechanism draws inspiration from Flash Attention [12]. We extend its block-wise streaming computation strategy to dynamic sparse scenarios by first partitioning $Q,K,V,Mask,Bias$ into blocks and sequentially loading them into global memory. If an entire block is masked, we skip the SRAM loading and attention weight computation for that block. For active blocks, we reuse the computation flow of Flash Attention. Finally, we concatenate the outputs block-wise and write them back to HBM. This design of skipping invalid computations block-wise combines the efficient memory access of Flash Attention with the content sparsity of DMA, reducing the total FLOPs from $O(n^{2}d_{h})$ to $O(nwd_{h})$ while fully utilizing the high throughput capabilities of GPU Tensor Cores.
This design produces the multi-level mask structure and weights structure shown in Figure 3: First, at the head level, unlike causal mask Self-Attention and semi-separable State-Space, Dynamic Mask Attention can focus on different contextual patterns in different heads. Its uniqueness lies in its ability to dynamically generate unique mask structures for each head based on content, allowing it to focus on different segments of information in different representation subspaces, maximizing the utilization of each subspace. Second, at the sequence position level, each position follows causal constraints and can only see past information. Finally, at the content level, it can both decay or enhance individual tokens by adding $\delta$ to the mask and dynamically determine which past positions are relevant by applying a top-k operation based on the values of $\delta$. This content-aware attention mask mechanism respects the sequential properties of language generation while improving computational efficiency and model expressiveness through dynamic allocation of attention weights, directly corresponding to the content sparsity in the selection task and associative sparsity in the perception task described earlier. It allows the model to adaptively focus on the most relevant contextual information while maintaining a total computational complexity of $O(n_{h}nwd_{h})$, with an upper limit of $n_{h}\times n\times w$ tokens that can be attended to.
Trainable Sparse Computing Optimization.
It is worth noting that in the actual computation process of Dynamic Mask Attention, the mask not only serves as a theoretical sparsity mechanism but also translates into significant computational efficiency improvements. When the mask value $m_{n_{h},j}=-\infty$, we can completely skip the QK product computation at the corresponding position. This optimization is not an approximation but mathematically precise, and it can significantly reduce the computational load while ensuring continuity of the gradient flow. Below, we provide a theoretical proof explaining why this skipping computation optimization is safe.
First, we analyze the forward propagation. For the masked position $(n_{h},j)$, its mask value $m_{n_{h},j}=-\infty$. When applying the softmax function, as shown in Equation 6, it results in 0, meaning that the attention weight $a_{n_{h},j}$ at that position must be 0. This implies that regardless of the result of the QK product computation, the attention weight at that position will not contribute to the final output. Therefore, we can safely skip the dot product computation for these positions and set the corresponding attention weights to 0, resulting in the same outcome as if we had performed the full computation followed by applying the mask. This optimization is particularly important in long sequence processing, as in modern inference scenarios, the length of the kv cache can reach thousands of tokens, and in some tasks, even millions of tokens. This makes directly computing the QK product for all positions extremely inefficient. By skipping the computation in this way, we can safely reduce the computation for $n-w$ positions and directly set the corresponding attention weights to 0, significantly improving computational efficiency.
$a_{n_{h},j}$ $=\frac{\exp(\frac{q_{n_{h}}\cdot k_{n_{h},j}^{\top}}{\sqrt{d_{h}}}+-\infty)}{\sum_{j^{\prime}=1}^{t}\exp(\frac{q_{n_{h}}\cdot k_{n_{h},j^{\prime}}^{\top}}{\sqrt{d_{h}}}+m_{n_{h},j^{\prime}})}=0$ (6)
Next, we analyze the backpropagation, which is crucial as it determines the training compatibility of the optimization. For backpropagation, we need to analyze the gradient flow. Let the loss function be $L$. For the query vector $q_{n_{h}}$ and key vector $k_{n_{h},j}^{\top}$ at the masked position $(n_{h},j)$, their gradients are computed using the chain rule, as shown in Equation 7.
$\frac{\partial L}{\partial q_{n_{h}}}$ $=\sum_{j^{\prime}=1}^{t}\frac{\partial L}{\partial a_{n_{h},j^{\prime}}}\frac{\partial a_{n_{h},j^{\prime}}}{\partial q_{n_{h}}}$ (7)
$\frac{\partial L}{\partial k_{n_{h},j}^{\top}}$ $=\frac{\partial L}{\partial a_{n_{h},j}}\frac{\partial a_{n_{h},j}}{\partial k_{n_{h},j}^{\top}}$
For masked positions, when $m_{n_{h},j}=-\infty$, we know that $a_{n_{h},j}=0$. More importantly, the derivative of the softmax function at this position with respect to its input is also zero: $\frac{\partial a_{n_{h},j}}{\partial s_{n_{h},j}}=0$, where $s_{n_{h},j}=\frac{q_{n_{h}}\cdot k_{n_{h},j}^{\top}}{\sqrt{d_{h}}}$. This is because the exponential of negative infinity in the softmax numerator results in zero, and the derivative of a constant zero with respect to any input is also zero. By applying the chain rule, we can derive the gradients for the query and key vectors at masked positions, as shown in Equation 8:
$\frac{\partial a_{n_{h},j}}{\partial q_{n_{h}}}$ $=\frac{\partial a_{n_{h},j}}{\partial s_{n_{h},j}}\frac{\partial s_{n_{h},j}}{\partial q_{n_{h}}}=0\cdot\frac{k_{n_{h},j}^{\top}}{\sqrt{d_{h}}}=0$ (8)
$\frac{\partial a_{n_{h},j}}{\partial k_{n_{h},j}^{\top}}$ $=\frac{\partial a_{n_{h},j}}{\partial s_{n_{h},j}}\frac{\partial s_{n_{h},j}}{\partial k_{n_{h},j}^{\top}}=0\cdot\frac{q_{n_{h}}}{\sqrt{d_{h}}}=0$
This analysis indicates that the key vector $k_{n_{h},j}^{\top}$ at masked positions does not receive any gradient through the masked attention weights. For the query vector $q_{n_{h}}$, it only receives gradient information through unmasked positions, as detailed in Equation 9.
$\frac{\partial L}{\partial q_{n_{h}}}$ $=\sum_{j^{\prime}=1}^{t}\frac{\partial L}{\partial a_{n_{h},j^{\prime}}}\frac{\partial a_{n_{h},j^{\prime}}}{\partial q_{n_{h}}}$ (9)
$=\sum_{j^{\prime}:m_{n_{h},j^{\prime}}\neq-\infty}\frac{\partial L}{\partial a_{n_{h},j^{\prime}}}\frac{\partial a_{n_{h},j^{\prime}}}{\partial q_{n_{h}}}+\sum_{j^{\prime}:m_{n_{h},j^{\prime}}=-\infty}\underbrace{\frac{\partial L}{\partial a_{n_{h},j^{\prime}}}\frac{\partial a_{n_{h},j^{\prime}}}{\partial q_{n_{h}}}}_{0}$
$=\sum_{j^{\prime}:m_{n_{h},j^{\prime}}\neq-\infty}\frac{\partial L}{\partial a_{n_{h},j^{\prime}}}\frac{\partial a_{n_{h},j^{\prime}}}{\partial q_{n_{h}}}$
Thus, through the above analysis, we have proven that skipping the QK computation at masked positions does not affect the model’s gradient flow. For $k_{n_{h},j}^{\top}$ at masked positions, its gradient should be 0; for the query vector $q_{n_{h}}$, it receives complete gradient information through unmasked positions. Therefore, we can safely skip these computations, significantly improving computational efficiency without affecting the model’s training process.
This theoretical analysis can guide us in developing specialized computation kernels that efficiently detect and skip masked regions at the CUDA level. Compared to traditional sparse computation methods, the sparsity optimization of DMA offers two key advantages: First, it maintains full differentiability during training, avoiding gradient issues caused by discretization operations; second, its block-level sparsity pattern is highly compatible with modern GPU architectures, effectively leveraging parallel computation and Tensor Core acceleration.
In conclusion, our trainable sparse computing optimization is not only theoretically precise and safe but also achieves significant performance improvements in practical implementations. In inference scenarios, each query can safely skip computations for up to $(n-w)$ positions. In training scenarios, it maintains the integrity of the gradient flow. DMA uses the same sparsification strategy in both training and inference phases, ensuring that the model can learn effective dynamic mask patterns during training and apply these patterns efficiently during inference. This design allows the model to learn end-to-end the optimal attention sparsity patterns, dynamically adjusting which historical positions are most critical for the current inference, thus achieving truly content-aware selective computation.
@@ 5.1 Experimental Settings
To thoroughly evaluate DMA, we benchmark it against the representative baselines surveyed in Section 2. First, we compared DMA with various mainstream attention variants in terms of pre-training perplexity across different model scales, and further validated DMA’s advantages in long-sequence information retrieval and computational efficiency in the challenging multi-query associative recall task, as well as its speed test in Section 5.2. Next, we tested the speedup of our kernel implementations compared to full attention in Section 5.3. Finally, we conducted a comparative evaluation of DMA, NSA, and MHA on a 1.7B parameter scale Transformer model pre-trained on 40B tokens, and assessed their performance on downstream benchmark tasks and needle-in-a-haystack tests in Section 5.4. These experiments demonstrate that DMA significantly improves computational efficiency while surpassing the performance of existing methods.
@@
All experiments were conducted using the open-source PyTorch images [39] and the Transformers framework [51]. For model configuration, we consistently employed the NeoX tokenizer [7], the AdamW optimizer [35], and the WSD learning rate scheduler [24], while strictly adhering to the Optimal Hyperparameter Scaling Law [30] and the Chinchilla [26] standard protocol throughout our training on the SmolLMCorpus [5] dataset. For evaluation frameworks, we utilized the LM evaluation harness [19] from EleutherAI for perplexity tasks, and the lighteval [17] from HuggingFace for downstream tasks.

### v6

@@ 4.2 Compute Position-Aware Sparse Weights
Algorithm 1 Flash Dynamic Mask Attention Forward Pass Per Head
0:  Matrices $\mathbf{Q},\mathbf{K},\mathbf{V}\in\mathbb{R}^{N\times d_{h}},\mathbf{M}\in\mathbb{R}^{N}$ in HBM. Set block sizes $B$.
  Initialize $\mathbf{O}=(0)_{N\times d_{h}}\in\mathbb{R}^{N\times d_{h}},\ell=(0)_{N}\in\mathbb{R}^{N},m=(-\infty)_{N}\in\mathbb{R}^{N}$ in HBM.
  Divide $\mathbf{Q}$ into $T_{r}$ blocks of size $B\times d_{h}$ each, and divide $\mathbf{K},\mathbf{V}$ into $T_{c}$ blocks of size $B\times d_{h}$ each.
  Divide $\mathbf{M}$ into $T_{c}$ blocks of size $B$ each.
  Divide $\mathbf{O}$ into $T_{r}$ blocks of size $B\times d_{h}$ each, divide $\ell$ into $T_{r}$ blocks of size $B$ each, divide $m$ into $T_{r}$ blocks of size $B$ each.
  for $1\leq j\leq T_{c}$ do
   Load $\mathbf{M}_{j}$ from HBM to SRAM.
   Compute $\mathrm{active}_{j}=\mathrm{Judge}(\mathbf{M}_{j})$.
   if $\mathrm{active}_{j}=0$ then
    Advance stream pointers for $\mathbf{K}_{j},\mathbf{V}_{j}$, continue.
   end if
   Load $\mathbf{K}_{j},\mathbf{V}_{j}$ from HBM to SRAM.
   for $1\leq i\leq T_{r}$ do
    Load $\mathbf{Q}_{i},\mathbf{O}_{i},\ell_{i},m_{i}$ from HBM to SRAM.
    Compute $\mathbf{S}_{ij}=\mathbf{Q}_{i}\mathbf{K}_{j}^{T}\times d_{h}^{-0.5}+\mathbf{M}_{j}\in\mathbb{R}^{B\times B}$.
    Compute $\tilde{m}_{ij}=\mathrm{rowmax}(\mathbf{S}_{ij})\in\mathbb{R}^{B}$, $\tilde{\mathbf{P}}_{ij}=\exp(\mathbf{S}_{ij}-\tilde{m}_{ij})\in\mathbb{R}^{B\times B}$, $\tilde{\ell}_{ij}=\mathrm{rowsum}(\tilde{\mathbf{P}}_{ij})\in\mathbb{R}^{B}$.
    Compute $m_{i}^{\mathrm{new}}=\max(m_{i},\tilde{m}_{ij})\in\mathbb{R}^{B}$, $\ell_{i}^{\mathrm{new}}=e^{m_{i}-m_{i}^{\mathrm{new}}}\ell_{i}+e^{\tilde{m}_{ij}-m_{i}^{\mathrm{new}}}\tilde{\ell}_{ij}\in\mathbb{R}^{B}$.
    Write $\mathbf{O}_{i}\leftarrow\mathrm{diag}(\ell_{i}^{\mathrm{new}})^{-1}(\mathrm{diag}(\ell_{i})e^{m_{i}-m_{i}^{\mathrm{new}}}\mathbf{O}_{i}+e^{\tilde{m}_{ij}-m_{i}^{\mathrm{new}}}\tilde{\mathbf{P}}_{ij}\mathbf{V}_{j})$ to HBM.
    Write $\ell_{i}\leftarrow\ell_{i}^{\mathrm{new}}$, $m_{i}\leftarrow m_{i}^{\mathrm{new}}$ to HBM.
   end for
  end for
  Return $\mathbf{O},\ell,m$.
This method offers three key advantages. First, the mask prunes the set of candidate tokens before the matrix multiplication and softmax operations. This avoids the inefficiency of pseudo-sparsity, where computations are performed for all tokens only to be zeroed out afterward. Second, unlike methods that perform key-value selection by discarding tokens, our approach preserves the full sequence. This ensures that the complete global context remains available for all attention heads to access as needed. Third, the kernel implementation can perform block-level skipping by loading a block of the mask to check if all positions within it are masked. If so, the entire block is skipped, avoiding unnecessary memory loads and matrix multiplication operations. The kernel implementation, depicted in Figure 2 (right), is optimized for this process. In an outer loop, blocks of the $K$ and $V$ matrices are loaded into SRAM. In an inner loop, blocks of the $Q$ matrix are loaded, and if the corresponding $K$ block is not entirely masked, the attention weights are computed and the output is written back to HBM. If a position in the $K$ block is masked, its attention weight is set to zero, and the computation for that position is skipped, resulting in position-aware sparse attention weights.
4.3 Fully Gradient Flow
Algorithm 2 Flash Dynamic Mask Attention Backward Pass Per Head
0:  Matrices $\mathbf{Q},\mathbf{K},\mathbf{V},\mathbf{O},\mathbf{dO}\in\mathbb{R}^{N\times d_{h}},\mathbf{M},\mathbf{dM}\in\mathbb{R}^{N}$ in HBM, vectors $\ell,m\in\mathbb{R}^{N}$ in HBM. Set block sizes $B$.
  Divide $\mathbf{Q}$ into $T_{r}$ blocks of size $B\times d_{h}$ each, and divide $\mathbf{K},\mathbf{V}$ in to $T_{c}$ blocks of size $B\times d_{h}$ each.
  Divide $\mathbf{M}$ into $T_{c}$ blocks of size $B$ each.
  Divide $\mathbf{O},\mathbf{dO}$ into $T_{r}$ blocks of size $B\times d_{h}$ each, divide $\ell$ into $T_{r}$ blocks of size $B$ each, and divide $m$ into $T_{r}$ blocks of size $B$ each.
  Initialize $\mathbf{dQ}=(0)_{N\times d_{h}}$ in HBM and divide it into $T_{r}$ blocks of size $B\times d_{h}$ each.
  Initialize $\mathbf{dK}=(0)_{N\times d_{h}},\mathbf{dV}=(0)_{N\times d_{h}}$ in HBM and divide $\mathbf{dK},\mathbf{dV}$ in to $T_{c}$ blocks of size $B\times d_{h}$ each.
  Initialize $\mathbf{dM}$ = (0)N in HBM and divide it into $T_{c}$ blocks of size $B$ each.
  for $1\leq j\leq T_{c}$ do
   Load $\mathbf{M}_{j}$ from HBM to SRAM.
   Compute $\mathrm{active}_{j}=\mathrm{Judge}(\mathbf{M}_{j})$.
   if $\mathrm{active}_{j}=0$ then
    Advance stream pointers for $\mathbf{K}_{j},\mathbf{V}_{j}$, continue.
   end if
   Load $\mathbf{K}_{j},\mathbf{V}_{j}$ from HBM to SRAM.
   Initialize $\tilde{\mathbf{dK}}_{j}=(0)_{B\times d_{h}},\tilde{\mathbf{dV}}_{j}=(0)_{B\times d_{h}},\tilde{\mathbf{dM}}_{j}=(0)_{B}$ on SRAM.
   for $1\leq i\leq T_{r}$ do
    Load $\mathbf{Q}_{i},\mathbf{O}_{i},\mathbf{dO}_{i},\mathbf{dQ}_{i},\ell_{i},m_{i}$ from HBM to SRAM.
    Compute $\mathbf{S}_{ij}=\mathbf{Q}_{i}\mathbf{K}_{j}^{T}\times d_{h}^{-0.5}+\mathbf{M}_{j}\in\mathbb{R}^{B\times B}$.
    Compute $\mathbf{P}_{ij}=\mathrm{diag}(\ell_{i})^{-1}\exp(\mathbf{S}_{ij}-m_{i})\in\mathbb{R}^{B\times B}$.
    Compute $\tilde{\mathbf{dV}_{j}}\leftarrow\tilde{\mathbf{dV}_{j}}+\mathbf{P}_{ij}^{\top}\mathbf{dO}_{i}\in\mathbb{R}^{B\times d_{h}}$.
    Compute $\mathbf{dP}_{ij}=\mathbf{dO}_{i}\mathbf{V}_{j}^{\top}\in\mathbb{R}^{B\times B}$.
    Compute $D_{i}=\mathrm{rowsum}(\mathbf{dO}_{i}\circ\mathbf{O}_{i})\in\mathbb{R}^{B}$.
    Compute $\mathbf{dS}_{ij}=\mathbf{P}_{ij}\circ(\mathbf{dP}_{ij}-D_{i})\in\mathbb{R}^{B\times B}$.
    Compute $\tilde{\mathbf{dM}}_{j}\leftarrow\tilde{\mathbf{dM}}_{j}+\mathrm{rowsum}(\mathbf{dS}_{ij})\in\mathbb{R}^{B}$.
    Write $\mathbf{dQ}_{i}\leftarrow\mathbf{dQ}_{i}+\mathbf{dS}_{ij}\mathbf{K}_{j}\times d_{h}^{-0.5}\in\mathbb{R}^{B\times d_{h}}$ to HBM.
    Compute $\tilde{\mathbf{dK}}_{j}\leftarrow\tilde{\mathbf{dK}}_{j}+\mathbf{dS}_{ij}^{\top}\mathbf{Q}_{i}\times d_{h}^{-0.5}\in\mathbb{R}^{B\times d_{h}}$.
   end for
   Write $\mathbf{dK}_{j}\leftarrow\tilde{\mathbf{dK}_{j}},\mathbf{dV}_{j}\leftarrow\tilde{\mathbf{dV}_{j}},\mathbf{dM}_{j}\leftarrow\mathbf{dM}_{j}+\tilde{\mathbf{dM}}_{j}$ to HBM.
  end for
  Return $\mathbf{dQ},\mathbf{dK},\mathbf{dV},\mathbf{dM}$.
Finally, the entire backward process is outlined in Algorithm 2. We ensure that the introduced dynamic mask and sparse weights do not block gradients, and the gradients of the retained attention paths are strictly consistent with those of full attention. They can flow completely to all inputs and parameters without gradient discontinuity issues caused by discrete operations, supporting end-to-end training while minimizing redundant costs.
For clarity, our derivation considers a single attention head $h$ at a single time step $t$. Let $\mathcal{I}_{h}\subset\{1,\dots,n\}$ be the set of $w$ indices selected for this head. For unselected indices $j\notin\mathcal{I}_{h}$, the mask value is treated as $m_{h,j}=$-$\infty$. The key intermediate variables in the forward pass are defined in Equation (6).
[…]
In the backward pass, let the upstream gradient of the loss function $L$ with respect to the head’s output $o_{h}$ be $do_{h}=\frac{\partial L}{\partial o_{h}}\in\mathbb{R}^{d_{h}}$. As shown in Equation (7), the gradient for $v$ is computed by distributing $do_{h}$ to the selected vectors $v_{h,j}$ in proportion to their attention weights $p_{h,j}$, while the gradients for unselected positions are zero.
$\frac{\partial L}{\partial v_{h,j}}=\begin{cases}p_{h,j}\,do_{h}&j\in\mathcal{I}_{h}\\ 0&j\notin\mathcal{I}_{h}\end{cases}$ (7)
Next, we compute the gradient for the scores $s_{h,j}$. The gradient of the attention weights $p_{h,j}$ with respect to their inputs is $dp_{h,j}=v_{h,j}\cdot do_{h}$. Using the standard softmax Jacobian, we can derive the gradient for $s_{h,j}$, denoted as $ds_{h,j}$, as shown in Equation (8). For masked positions where $p_{h,j}=0$, the gradient $ds_{h,j}$ is naturally zero.
$ds_{h,j}=p_{h,j}(dp_{h,j}-\sum_{j^{\prime}\in\mathcal{I}_{h}}p_{h,j^{\prime}}\times dp_{h,j^{\prime}})$ (8)
Because $s_{h,j}$ is an additive combination of $q_{h}\cdot k_{h,j}$ and $m_{h,j}$, the gradient is distributed directly. The gradient for the mask $m_{h,j}$ is simply $ds_{h,j}$, as shown in Equation (9). This ensures that gradients can flow directly to $\Delta$ and $A$.
$\frac{\partial L}{\partial m_{h,j}}=ds_{h,j}$ (9)
Finally, as shown in Equation (10), the gradients for $q_{h}$ and $k_{h,j}$ are obtained by backpropagating $ds_{h,j}$ through the computation path. Crucially, the gradient calculations only involve the selected index set $\mathcal{I}_{h}$, thereby reducing computation.
$\frac{\partial L}{\partial q_{h}}=\sum_{j\in\mathcal{I}_{h}}ds_{h,j}\frac{k_{h,j}}{\sqrt{d_{h}}},\qquad\frac{\partial L}{\partial k_{h,j}}=ds_{h,j}\frac{q_{h}}{\sqrt{d_{h}}}$ (10)
Our approach has several significant advantages. First, for the selected positions, the gradients are identical to those of full attention, and DMA only prunes the operator chain for positions whose contributions can be ignored, ensuring expressiveness. Then, only second-order correlation information is propagated to $\mathcal{I}_{h}$, improving bandwidth utilization. The gating parameter $A$ and weight $\Delta$ directly receive attention weights as gradients, quickly shaping head specialization. Finally, the equivalence relation $dM=dS$ allows the kernel to only recompute the local $S$ without storing additional intermediate mask gradient tensors.
@@ 5.1 Experimental Settings
To thoroughly evaluate DMA, we benchmarked it against representative baselines surveyed in Section 2. First, we compared DMA with various mainstream attention variants in terms of pre-training perplexity at different model scales, further validating DMA’s advantage in long-sequence information retrieval through the challenging multi-query associative recall task, and tested the efficiency of our kernel compared to other optimized implementations in Section 5.2. Second, we compared DMA, NSA, and MHA on a 1.7B parameter Transformer model pre-trained and supervisedly fine-tuned on 40B tokens, evaluating them on downstream benchmark tasks and needle-in-a-haystack tests in Section 5.3.
@@
All experiments were conducted using the open-source PyTorch images [nv2022pytorch] and the Transformers framework [wolf-etal-2020-transformers]. For model configuration, we consistently employed the NeoX tokenizer [black2022gpt], the AdamW optimizer [Loshchilov2017FixingWD], and the WSD learning rate scheduler [hägele2024scalinglawscomputeoptimaltraining], while strictly adhering to the Optimal Hyperparameter Scaling Law [li2025predictablescalei] and the Chinchilla [hoffmann2022empirical] standard protocol throughout our training on the SmolLMCorpus [benallal2024smollmcorpus] dataset. For evaluation frameworks, we utilized the LM evaluation harness [eval-harness] from EleutherAI for perplexity tasks, and the lighteval [Clémentine2023lighteval] from HuggingFace for downstream tasks.